1/2

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,
Varulven  The Werewolf  The Banshee (An Approach)
tradukita de Sven Collberg  tradukita de Alexander Gross  tradukita de Max Knight

En varulv lopp från fru m.m. en vacker afton samt begav sig till en bymagisters grav och bad: “Var snäll och konjugera.

A Werewolf, troubled by his name, Left wife and brood one night and came To a hidden graveyard to enlist The aid of a long-dead philologist.

One night, a banshee slunk away

from mate and child, and in the gloom went to a village teacher’s tomb, requesting him: “Inflect me, pray.”

mig!” Byskolläraren stod upp på blecknamnsskyltens mässingknopp med svar till den, som satt på pass med tåligt korslagd tass på tass.

“Oh sage, wake up, please don’t berate me,”

He howled sadly, “Just conjugate me.”

The seer arose a bit unsteady

Yawned twice, wheezed once, and then was ready.

The village teacher climbed up straight upon his grave stone with its plate and to the apparition said who meekly knelt before the dead:

“Jag var-ulv 1:a pers. i sing., du, han, hon, den, det var-ulv. Så vi voro-ulv, I voren. På de voro-ulv kom ingenting.”

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

“The banSHEE, in the subject’s place; the banHERS, the possessive case. The banHER, next, is what they call objective case–and that is all.”

Varulven myste gladåt hur pers. hade följts av sing. och plur. “Fast”, bad han, “kan det inte ges ens en enda pers. i tempus presens?”

“I know that–I’m no mental cripple–

The future form and participle

Are what I crave,” the beast replied. The scholar paused–again he tried:

The banshee marveled at the cases

and writhed with pleasure, making faces, but said: “You did not add, so far, the plural to the singular!”

Bekänna måste vår magister, att därvidlag fanns vissa brister. En varulv sine imperfecto var contradictio in adjecto.

“A ‘Will-be-wolf?’ It’s just too long:

‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly wrong! Such words are wounds beyond all suture– I’m sorry, but you have no future.”

The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advised, “A she cannot be plurized.”

…  …  …

2/2

Varulven satt med tårad lins. Han visste att han fanns och finns. Därtill kom också fru m.m. som även måste existera.

The Werewolf knew better–his sons still slept At home, and homewards now he crept, Happy, humble, without apology For such folly of philology.

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the teacher.

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Svedan de Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

http://www.poezio.net/version?poem-id=825&version-id=1654,1657,1660